A clothes store buys a shirt for $20, and adds on 20% of this amount for its selling price. At a sale four weeks later, the customer is invited to take 20% off the price of the shirt. Does the store gain money, lose money,or break even when it sells the shirt on sale?
Answer by checkley75(3666)

20+.2*20=20+4=24 PRICE OF THE SHIRT BEFORE SALE
24-(.2*24)=24-4.8=19.6 PRICE OF SHIRT AFTER DISCOUNT.
THUS THEY LOOSE $.40 ON EACH SHIRT ON SALE.